# A Mixture-of-Experts Framework with Log-Logistic Components for Survival Analysis on Histopathology Images

### v2

@@ 5.2 Ablation Studies and Interpretability
Ablation studies are conducted to evaluate the contributions of Quantile-Gated Patch Selection, Graph-Guided Clustering, Hierarchical Context Attention and Expert-Driven Mixture of Log-logistics(Table 1(b)). Varying the quantile parameter showed that q=0.25q{=}0.25 provided the best trade-off, preserving prognostic regions while filtering irrelevant tissue. Eliminating any module resulted in a reduction in TDC, confirming their contributions to noise suppression and contextual feature integration. Single-expert model fell short, whereas five experts yielded superior results, underscoring the benefit of multi-expert modeling. For interpretability, patients were categorized into high and low-risk groups based on their predicted survival scores. The Kaplan–Meier curves(Fig. 2) revealed distinct survival trends with significant separation (log-rank p=0.046p{=}0.046 for LUAD, p=0.005p{=}0.005 for KIRC, p=0.017p{=}0.017 for BRCA), confirming that the expert-driven architecture effectively captures prognostically relevant morphological patterns.
@@ 6 Conclusion
We proposed a unified framework for WSI-based survival prediction integrating quantile-based patch selection, graph-guided clustering, hierarchical context attention, and expert-driven mixture of log-logistics. By modeling local–global tissue interactions through clustered attention and estimating survival distributions via multiple experts, the method improves calibration and discrimination. Experiments on TCGA-LUAD, TCGA-KIRC, and TCGA-BRCA show consistent gains in time-dependent concordance over pathology-based and multimodal baselines. Future work will extend the framework towards multimodal and uncertainty-aware patient-wise survival prediction for clinical reliability.

### v3

@@ 4.2 Ablation Studies and Component Analysis
Ablation studies evaluate the contributions of Quantile-Gated Patch Selection, Graph-Guided Clustering, Hierarchical Context Attention and Expert-Driven Mixture of Log-logistics (Table 1b). Varying the quantile parameter shows that $q=0.25$ provides the best trade-off, preserving prognostic regions while filtering irrelevant tissue. Eliminating any module results in reduced TDC, confirming their complementary contributions to noise suppression and contextual feature integration. The single-expert configuration ($e=1$) underperforms compared to five experts, underscoring the benefit of multi-expert modeling.
4.3 Clinical Validation and Interpretability
For interpretability, patients are categorized into high- and low-risk groups based on predicted survival scores. The Kaplan-Meier curves (Fig. 2) reveal distinct survival trends with statistically significant separation (log-rank $p=0.046$ for LUAD, $p=0.005$ for KIRC, $p=0.017$ for BRCA), confirming that the proposed method effectively captures prognostically relevant morphological patterns. The clear risk stratification demonstrates clinical utility for identifying patients who may benefit from more aggressive treatment.
@@ 5 Conclusion
We presented a unified computational pathology framework for WSI-based survival prediction that integrates quantile-based patch selection, graph-guided clustering, hierarchical context attention, and expert-driven mixture modeling. Our approach captures local–global tissue interactions through spatially coherent clustering and models complex survival distributions via specialized log-logistic experts, achieving both superior discrimination and calibration. Extensive validation on TCGA-LUAD, KIRC, and BRCA cohorts demonstrates consistent and statistically significant improvements, outperforming both pathology-based and multimodal benchmarks. Notably, our framework surpasses multimodal methods using histology alone, highlighting untapped prognostic value in morphological patterns. Future work will explore multimodal integration and uncertainty estimation for greater clinical robustness.